OpenStudy (anonymous):

find the surface are of the surface generated by revolving the curve x=3/8(y4/3−2y2/3) from y=1 to y=8 revolving about the x-axis. so far this is what i have Is any of this correct? ∫183/8(y4/3−2y2/3) 3/8[3/7y7/3−2(3/5)5/3]from8→1 3/8[(3/7)(8)7/3−2(3/5)(8)5/3]−3/8[(3/7)(1)7/3−2(2/5)(1)5/3] (3/8)[(387/7)−(192/5)]−(3/8)[(3/7)−(6/5)] (1773/280)−(291/280) =740/140

OpenStudy (anonymous):

its correct so farrr

OpenStudy (anonymous):

For surface area it is: \[2 \pi \int\limits_a^b r(y)\sqrt{1+(\frac{dx}{dy})^2}dy\]

OpenStudy (anonymous):

In this case r(y)=y